OpenStudy (anonymous):

Choose the correct description of the graph of the inequality −x greater than or equal to 3. Open circle on −3, shading to the left. Closed circle on −3, shading to the left. Open circle on 3, shading to the right. Closed circle on 3, shading to the right.

terenzreignz (terenzreignz):

Someone misses inequalities... anyway, @dmezzullo the words say "-x greater than or equal to 3" :P

OpenStudy (dmezzullo):

lol wait was i right?

terenzreignz (terenzreignz):

Uhh... I don't know... \[\large -x \ge 3\]

OpenStudy (dmezzullo):

@bella0113

terenzreignz (terenzreignz):

bolted.

jimthompson5910 (jim_thompson5910):

\[\large -x \ge 3\] would turn into \[\large x \le -3\] because changing the sign of each term flips the inequality sign